Given the function f(x) = 4(x+3) − 5, solve for the inverse function when x = 3.
Answer by CubeyThePenguin(3113)

Find the inverse first.

y = 4(x+3)- 5
y + 5 = 4(x + 3)
(y + 5)/4 = x + 3
(y + 5)/4 - 3 = x

f^(-1)(x) = (x + 5)/4 - 3

f^(-1)(3) = (3 + 5)/4 - 3 = -1